OpenStudy (anonymous):

how can i solve a quadratic equation like this 6x^2-5+x=0

OpenStudy (anonymous):

rewrite it in the normal form and plug into the quadratic equation: \[ax^2 + bx + c = 0 \implies x = \frac{-(b) \pm \sqrt{(b)^2-4(a)(c)}}{2(a)}\]

OpenStudy (anonymous):

thank you i did that part but where i am getting cofused at is that addition/ subtraction sign i dont kno what to do with it sorry i didnt say that earlier

OpenStudy (anonymous):

Oh, well unless the part under the square root is 0 you will get 2 different possible solutions. \[x = \frac{-(b) + \sqrt{(b)^2-4(a)(c)}}{2(a)}\] and \[x = \frac{-(b) - \sqrt{(b)^2-4(a)(c)}}{2(a)}\] The \(\pm\) thing just makes it easier to write.

OpenStudy (anonymous):

oh ok thank you i'ma try it now :D

OpenStudy (anonymous):

HOW WOUD I GET THAT SQUARE ROOT SYMBOL ON THE CALCULATOR

OpenStudy (anonymous):

It should have it. If not you can raise it to the .5 power.

OpenStudy (anonymous):

LIKE THIS (1^2-4(6)(-5)^5

OpenStudy (anonymous):

it gave me some crazy number how would you do it i guess i was wrong

OpenStudy (anonymous):

i got 2.59374246E10

OpenStudy (anonymous):

Not quite. What do you have before you plug it in?

OpenStudy (anonymous):

i did 1^2-4(6)(-5) and i recieved 121 but then i raised that to the 5th power and i got that wierd number

OpenStudy (anonymous):

but then instead of that i subtracted -5 then i divided by 12 and i got 10.5

OpenStudy (anonymous):

?

OpenStudy (anonymous):

why did you raise to the 5th. I said .5 or 0.5 or 1/2

OpenStudy (anonymous):

that's probably the issue.

OpenStudy (anonymous):

i still dnt get it i'm sorry for wasting your time!

OpenStudy (anonymous):

Your calculator doesn't have a square root button?

OpenStudy (anonymous):

it does but it is a blue sign above the X^2 button i cnt seem to get the sign but i always get the X^2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

What kind of calculator do you have

OpenStudy (anonymous):

a TI-84 PLUS SILVER ADDITION

OpenStudy (anonymous):

Do you see the blue func button? Press that first, then press the x^2 button.

OpenStudy (anonymous):

err should say 2nd on it.

OpenStudy (anonymous):

I GOT IT! yah thanx yu

OpenStudy (anonymous):

i got a -48

OpenStudy (anonymous):

The square root of 121 is 11

OpenStudy (anonymous):

So you should have \[x = \frac{-1 \pm 11}{2(6)}\]

OpenStudy (anonymous):

i feel so dum sorry!

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

hrm?

OpenStudy (anonymous):

i got 32and -34 do i put down both of my answers

OpenStudy (anonymous):

How do you get that?

OpenStudy (anonymous):

-1 + 11 = 10 -1 - 11 = -12 So you should have 10/12 and -12/12 = -1

OpenStudy (anonymous):

omg! im sorry but im dne i cnt do this!

OpenStudy (anonymous):

I don't understand what's confusing you.

OpenStudy (anonymous):

If you can explain I can try to help.

OpenStudy (anonymous):

i just dont know how to do this im basically confused about evrything i dnt know what else to do

OpenStudy (anonymous):

Well hang on.

OpenStudy (anonymous):

Can you simplify the expression \[\frac{4+3}{8}\]

OpenStudy (anonymous):

7/8

OpenStudy (anonymous):

1.142857143?

OpenStudy (anonymous):

no im sorry 7/8

OpenStudy (anonymous):

u k i have no idea wat ur sayin im in 5th grade '

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Yes, 7/8 is correct

OpenStudy (anonymous):

Now try this one: \[\frac{4-3}{2(4)}\]

OpenStudy (anonymous):

polpok ur awesome

OpenStudy (anonymous):

heh, thanks.

OpenStudy (anonymous):

np

OpenStudy (anonymous):

When working with fractions like this, just try to simplify the top and bottom separately.